A total of 150 tickets were sold for the annual concert to students and non students. Student tickets were $5 and non-students tickets were $8. If the total revenue for the concert was $930, then how many tickets of each type were sold?
Answer by jorel1380(3719)

5p+8(150-p)=930
5p+1200-8p=930
3p=270
p=90
150-p=60
There were 90 student tickets sold and 60 non-student..